For ThF4 ksp= 5.0x10^-29. What is the solubility of the sale in water? Show all work
Answer #1

ThF4----> Th+4 + 4F-

Ksp = [ Th+4] {F-]4

Let [Th+4 ] =x hence [F-] =4x

5*10-29= x(4x)4= 256x5

x5= 5*10-29/ 256

x5= 1.95*10-31

x= 7.21*10-7M